OpenStudy (anonymous):

please help? i'm stuck. what is the x coordinate of the solution to the system? 2x-y-z=11 x+y-4z=-1 -x+3y+z=-10

OpenStudy (texaschic101):

x + y - 4z = -1 -x + 3y + z = -10 ----------------add 4y - 3z = -11 2x - y - z = 11 x + y - 4z = -1 -- multiply by -2 ------------- 2x - y - z = 11 -2x - 2y + 8z = 2 -------------add -3y + 7z = 13 4y - 3z = -11 --multiply by 3 -3y + 7z = 13 -- multiply by 4 ------------ 12y - 9z = -33 -12y + 28z = 52 ------------add 19z = 19 z = 1 -3y + 7z = 13 -3y + 7(1) = 13 -3y + 7 = 13 -3y = 13 - 7 -3y = 6 y = -6/3 y = -2 x + y - 4z = -1 x + (-2) - 4(1) = -1 x - 2 - 4 = -1 x - 6 = -1 x = -1 + 6 x = 5 check... 2x - y - z = 11 2(5) - (-2) - 1 = 11 10 + 2 - 1 = 11 12 - 1 = 11 11 = 11 (correct) x = 5, y = -2 and z = 1